OpenStudy (anonymous):

A right triangle has a hypotenuse of sqrt(18) inches. Write a function A(x) that represents the area of the triangle

OpenStudy (precal):

A=(bh)/2

OpenStudy (precal):

a^2+b^c=18^2

OpenStudy (precal):



OpenStudy (precal):

I would take a^2+b^2=18^2 and solve for a a=square root (18^2-b^2)

OpenStudy (precal):